### v1

# The Jaynes Cummings model as an autonomous Maxwell demon

Abstract
We revisit the Jaynes cummings model as an autonomous thermodynamic machine, where a qubit is coupled to a cavity with a very large initial coherent field. Our analysis reveals different behaviors of this model including the expected unitary behavior of qubit at short times, which fails at longer times when the cavity starts measuring the qubit. The subsequent dynamics feature feedback from the cavity on the qubit, which appears to purify the qubit irrespective of its initial state. Upon analyzing these dynamics from a thermodynamic standpoint, we found that the cavity behaves as an autonomous Maxwell demon.
Autonomous machines, performing tasks without external control, are an important concept in thermodynamics [1, 2]. Their time-dependent evolution derive solely from their initial non-equilibrium condition, akin to automaton. This concept is particularly important in the case of quantum machines [3, 4, 5, 6, 7, 8, 9], where the coherence of microscopic constituents is crucial, as avoiding classical controllers empirically mitigates decoherence. This motivates research into autonomous (or coherent) feedback [10, 11, 12, 13, 14, 15], where a quantum system is able to drive a target system conditionally on the values of a few of its observables, e.g. applied to quantum error correction [16, 17, 18]. When these controllers perform thermodynamic tasks, such as work extraction or cooling, they are called autonomous Maxwell demons [19, 20, 21, 22, 23, 24, 25]. A thorough characterization of these machines is of major fundamental interest, as it is linked to the long-standing question of how the dynamics induced by quantum measurements [26, 27, 28] and the laws of thermodynamics [29, 30, 31] emerge in an entirely quantum framework.
Here, we analyze an autonomous machine composed of a qubit interacting with a cavity initialized in a coherently displaced thermal state, focusing on the classical limit. Over time, we identify three successive regimes exhibiting very different thermodynamic behaviors. At short times, the cavity behaves as a quasi-ideal source of work, namely, a classical drive inducing unitary Rabi oscillations of the qubit. We quantify the deviations from ideality, which take the form of residual heat exchange and entropy production. This regime breaks down over a time-scale set by the qubit-cavity coupling constant, where Rabi oscillations decohere. This decoherence is related to an autonomous measurement of the qubit by the cavity in a basis set by the initial phase of the field. The result of this measurement is stored in a conditional field amplitude. Finally, we identify a third regime where the cavity unitarily drives the qubit depending on these measurement outcomes. This autonomous feedback is able to purify the qubit starting from any mixed initial state. Our thermodynamic analysis shows that the cavity behaves as an autonomous Maxwell demon, trading mutual information for cooling power.
Model – We consider qubit $Q$ (e.g. a two-level atom) interacting with a harmonic oscillator (hereafter dubbed the cavity $C$) according to the Jaynes-Cummings interaction $\hat{V}_{\text{QC}}=i\hbar g(a\sigma_{+}-a^{\dagger}\sigma_{-})$, where $a$ denotes the bosonic annihilation operator of the cavity, and $\sigma_{-}=|g\rangle\langle e|=\sigma_{+}^{\dagger}$ the lowering operator of the atom. We focus on resonant conditions, where the qubit and cavity Hamiltonians are $\hat{H}_{Q}=\hbar\omega_{0}\sigma_{+}\sigma_{-}$ and $\hat{H}_{C}=\hbar\omega_{0}a^{\dagger}a$, respectively. We assume that the cavity is initially prepared in a coherently-displaced thermal state $\hat{\rho}_{C}(0)=D(\alpha_{0})\hat{w}_{\beta(0)}D^{\dagger}(\alpha_{0})$, where $\hat{w}_{\beta(0)}=\frac{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}}{Z_{C}(0)}$ is a thermal equilibrium state at inverse temperature $\beta(0)$, with $Z_{C}(0)=\text{Tr}\{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}\}$. $D(\alpha_{0})=e^{\alpha_{0}a^{\dagger}-\alpha_{0}^{*}a}$ denotes the phase-space displacement operator with amplitude $\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}$. We are interested in the limit $n_{0}\gg\bar{n},1$, with $\bar{n}=(e^{\beta(0)\hbar\omega_{0}}-1)^{-1}$ the initial thermal occupation of the cavity, in which the cavity is expected to induce a classical drive for the qubit. Due to the resonance condition, the evolution of the total system in the interaction picture is governed by Hamiltonian $\hat{H}_{I}(t)=V_{\text{QC}}$ [32].
Figure 1: (a) Entropy variation $\Delta S_{Q}$ of the qubit over a long time scale for different initial qubit states $|\psi_{1}\rangle=|e\rangle\langle e|$, $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+e^{i\frac{\pi}{4}}|g\rangle)$, and $|\psi_{3}\rangle=|+_{y}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+i|g\rangle)$. (b) zoom in the early times of (a), in semilog scale.(c) Excited state population $P_{e}=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ (in blue) and coherence $C_{eg}=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ (in orange and red) of the atom when it is initialized in $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$. The dots correspond to expansions up to second order in $1/n_{0}$ and the solid lines are obtained from exact expressions, truncating the cavity Hilbert space to $N_{\text{ph}}=500$. Parameters: $n_{0}=100$, $\bar{n}=1$.
[…]
The autonomous dynamics of the machine features three successive regimes with very different thermodynamic interpretations, that can be identified from the time-evolution of the qubit’s von Neumann entropy as shown in Fig. 1a) and b). We analyze them in the following.
Unitary regime – For times $t\ll g^{-1}$, the cavity and the qubit density operators remain approximately factorized, such that the cavity behaves as a quasi-ideal source of work inducing a classical drive on the qubit. As a consequence, the qubit follows a unitary evolution ruled by the effective Hamiltonian $\hat{H}^{\text{eff}}_{Q}(t)=\text{Tr}_{C}\{\hat{V}_{\text{QC}}\hat{\rho}_{C}(t)\}=i\hbar g(\left\langle a(t)\right\rangle\sigma_{+}-\langle{a^{\dagger}(t)}\rangle\sigma_{-})$, featuring Rabi oscillations with frequency $\Omega\simeq g|\left\langle a(0)\right\rangle|=g\sqrt{n_{0}}$ (see Fig. 1(c) and [32]). To identify thermodynamic evidence of the work source behavior of the cavity, and deviations from it, we plot in Fig. 2 the work $W_{C}$ and the heat $Q_{C}$ provided by the cavity to the qubit, together with the variation of the cavity’s internal energy $\Delta E_{C}$. We see that for large enough values of $n_{0}$, the cavity mostly provides work and almost no entropy is produced during a single Rabi oscillation. We formalize this result by expanding the exact cavity and qubit state $\hat{\rho}(t)=\hat{U}(t)(\hat{\rho}_{Q}(0)\otimes\hat{\rho}_{C}(0))\hat{U}^{\dagger}(t)$ as power series of $1/n_{0}$. For $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, the qubit population $P_{e}(t)=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ and coherence $C_{eg}(t)=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ read:
$P_{e}(t)$ $=$ $\cos^{2}{\frac{\theta}{2}}+\frac{\delta P_{e}(\theta,\bar{n})}{n_{0}}+{\cal O}\!\left(\!\frac{1}{n_{0}^{2}}\!\right)$
$C_{eg}(t)$ $=$ $-e^{i\phi_{0}}\left[\frac{1}{2}\sin{\theta}+\frac{\delta C_{eg}(\theta,\bar{n})}{n_{0}}\right]+{\cal O}\!\left(\!\frac{1}{n_{0}^{2}}\!\right)\!\!,\quad\;\;$ (1)
where $\theta=2\Omega t=2gt\sqrt{n_{0}}$ stands for the rotation angle in the Bloch sphere. The complete expression of first order corrections $\delta P_{e}(\theta,\bar{n})/n_{0}$ and $\delta C_{eg}(\theta,\bar{n})/n_{0}$ to an ideal Rabi oscillation are given in the Supp. Mat., and decay as $1/n_{0}$ in the classical limit. These first-order truncations show good agreement with exact numerical results over short time-scales of the order of $\Omega^{-1}$, as shown in Fig. 1(c). In addition, our expansion shows that the amount of exchanged heat $Q_{C}$ and the entropy production scale as $1/n_{0}$ (see inset of Fig. 2), confirming that the cavity can be considered as an ideal work source in this regime.
Figure 2: Heat $Q_{C}$ (red), work $W_{C}$ (green) and internal energy variation of the cavity $\Delta E_{C}$ (blue) as a function of $\theta=2gt\sqrt{n_{0}}$ for $n_{0}=500$ and $\bar{n}=1$, obtained from the second order in $1/n_{0}$. Exact expressions with truncated cavity Hilbert space are displayed in black (not distinguishable). Inset: Heat provided by the cavity during a single Rabi oscillation ($\theta=2\pi$), obtained from the expansion, as a function of $n_{0}$, in semilog scale.
Measurement-induced decoherence – It is known from previous calculations [33], as well as measurements on Rydberg atoms in cavities [34], that the qubit decoheres, and Rabi oscillations collapse, over a time scale $t_{c}\sim g^{-1}$ (see Fig. 1c). This decoherence is due to the evolution of the field towards two orthogonal states, with macroscopically different values of the complex field amplitudes $\left\langle a(t)\right\rangle$ conditioned on the state of the qubit. This evolution can be interpreted as a measurement of the qubit by the cavity  [^1] A more correct terminology here would be a “pre-measurement”. The measurement would be complete only if the measurement outcome is copied in a large number of environmental degrees of freedom [35]. If the cavity and qubit continue to evolve in isolation, revivals of the Rabi oscillations would eventually be observed [33]. However, the measurement analogy remains fully consistent on the time interval considered in this letter.[36].
The measurement basis is set by the eigenbasis of $\hat{H}_{Q}^{\text{eff}}(0)$, and therefore by the initial phase of the field. Explicitly, it reads $\{\cos(\frac{\phi_{0}}{2})|+_{y}\rangle-i\sin(\frac{\phi_{0}}{2})|-_{y}\rangle,-i\sin(\frac{\phi_{0}}{2})|+_{y}\rangle+\cos(\frac{\phi_{0}}{2})|-_{y}\rangle\}$, where |±$|\pm_{y}\rangle$ are the eigenstates of $\sigma_{y}=i(\sigma_{-}-\sigma_{+})$, and corresponds to an axis in the equator of the Bloch sphere. In the case $\phi_{0}=0$, which we consider hereafter, the qubit is measured along the $y$-axis. The total state verifies for $t\lesssim t_{c}={\cal O}(g^{-1})$ [36]:
$\hat{\rho}(t)=\!\!\!\sum_{\nu,\mu=\pm}\rho_{\nu\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{C}^{\nu\mu}(t)+{\cal O}\!\left(\frac{1}{n_{0}}\right)\!,\;\;\;$ (2)
where $\rho_{\nu\mu}=\langle\nu_{y}|\rho_{Q}(0)|\mu_{y}\rangle$ are the coefficients of the qubit density operator in the measurement basis and
$\hat{\rho}_{C}^{\nu\mu}(t)\!\!\!$ $=$ $\!\!\!e^{\nu i\chi\delta_{\nu\mu}}D(\sqrt{n_{0}}e^{\nu\frac{igt}{2\sqrt{n_{0}}}})w_{\beta(0)}D^{\dagger}(\sqrt{n_{0}}e^{\mu\frac{igt}{2\sqrt{n_{0}}}}),\;\;\;$ (3)
are conditional operators in the cavity space with $\chi=gt(\sqrt{n_{0}}+1/\sqrt{n_{0}})$ and $\delta_{\nu\mu}$ the Kronecker delta. The diagonal terms $\hat{\rho}_{C}^{\nu\nu}$ are conditional cavity states associated with measurement results $\nu$, and are characterized by field amplitudes
$\alpha_{\nu}(t)=\sqrt{n_{0}}e^{\nu\frac{igt}{2\sqrt{n_{0}}}}.$ (4)
For $t\in[0,t_{c}]$, they fulfill $\alpha_{\nu}(t)=\sqrt{n_{0}}+i\nu\frac{gt}{2}+{\cal O}\left(1/\sqrt{n_{0}}\right)$. The trace of the cross terms $\text{Tr}\{\hat{\rho}_{C}^{+-}(t_{c})\}$ vanishes for $n_{0}\gg\bar{n},1$, which implies maximal qubit-cavity correlations, and hence a complete measurement [37]. In particular, the reduced qubit state verifies $\hat{\rho}_{Q}(t_{c})=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+{\cal O}(1/n_{0})$, achieving complete dephasing of the Rabi oscillations in the classical limit.
Autonomous feedback and qubit purification. Remarkably, the behavior of the machine presents a third regime on an even longer timescale $t\sim\sqrt{n_{0}}g^{-1}$, in which the cavity performs an effective drive depending on the result of the measurement in the {|±$\{|\pm_{y}\rangle\}$ basis, thereby performing an autonomous feedback. We show below that this feedback is able to purify any initial qubit state when $n_{0}$ is sufficiently large. We first note that at $t\sim t_{c}$, the cavity field $\text{Tr}\{a\hat{\rho}_{C}^{\nu\nu}\}=\alpha_{\nu}(t)$ depends on the measurement result $\nu$. It is instructive to analyze the joint qubit-cavity dynamics along the two independent branches associated with the qubit found in |±$|\pm_{y}\rangle$. As, inside each branch, the qubit-cavity state is factorized up to ${\cal O}(1/n_{0})$ corrections, one can again consider the cavity as an ideal work source, inducing a classical drive on the qubit of the form $\hat{H}^{\text{eff}(\nu)}_{Q}(t)=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-})$ [32], where, unlike early times ($t<t_{c}$), the effective driving Hamiltonian depends on the measurement result. At $t_{c}$, the two Hamiltonians almost coincide with $\hat{H}_{Q}^{\text{eff}}(0)$, and the qubit measured state is aligned with the field, whatever $\nu$. Then, two different field axes emerge. From Eq. (4), we see that they both lie in the equator and rotate in opposite directions with angular velocities ±$\pm\frac{g}{2\sqrt{n_{0}}}$ much smaller, by a factor of $1/n_{0}$, than the energy splitting of $\hat{H}^{\text{eff}(\nu)}_{Q}(t)$, still given by the Rabi frequency $\Omega$. As a consequence, the qubit follows adiabatically the rotation axis inside each measurement branch (see Fig. 3). The two field axes meet again at $t=t_{\text{min}}=\sqrt{n_{0}}\pi g^{-1}$ where they overlap with the $x$-axis of the Bloch sphere, driving in both branches the qubit to state $|+_{x}\rangle$. As a consequence of this autonomous feedback, an arbitrary initial qubit state $\hat{\rho}_{Q}(0)$ is deterministically purified and brought to $|+_{x}\rangle$ (or generally for a phase $\phi_{0}$ to state $(1/\sqrt{2})\left(|g\rangle+e^{i\phi_{0}}|e\rangle\right)$) at time $t_{\text{min}}$.
Figure 3: Autonomous feedback mechanism. (a) Trajectory of the qubit on the equator of the Bloch sphere starting from states $|+_{y}\rangle$ (blue) and $|-_{y}\rangle$ (green) and (b) of the average field amplitude. The red circle locates the qubit state reached at $t_{\text{min}}$. The yellow circle locates the initial cavity displacement, for $\phi_{0}=0$.
It is worth noting that this ideal purification occurs in the limit $n_{0}\to\infty$, such as $t_{\text{min}}$ also diverges. Nevertheless, our expansion at first order in ${\cal O}(1/n_{0})$ allows us to identify corrections for finite $n_{0}$. Mainly, the corrections are due to correlations between the cavity and the qubit inside each branch, which are responsible for a purity loss of the qubit state. As a result, the qubit achieves a non-zero minimum entropy (scaling as ${\cal O}(1/n_{0})$), at a time $t_{\text{min}}$ which slightly differs from $\sqrt{n_{0}}\pi$ (See Fig. 1(a)). In summary, there is a trade-off between the purity of the attained qubit state and the time needed to purify the qubit.
As seen in Fig. 1(a), the entropy of the qubit goes through a series of minima with a decreasing purity at times $t_{k}\simeq(2k+1)\sqrt{n_{0}}\pi$ (with $t_{1}\equiv t_{\text{min}}$), except when it starts already in one of the two pointer states |±$|\pm_{y}\rangle$ (red curve). In this case, the cavity-qubit correlations associated with the measurement do not build up, and the qubit entropy increase is only due to intra-branch correlations associated with finite $n_{0}$ behavior. Between two consecutive minima, the qubit goes through entropy maxima of increasing value. After a few oscillations, as intra-branch qubit-cavity correlations keep growing, the qubit entropy saturates at $\log 2$, the maximum physical value for a qubit.
Thermodynamic signature of an autonomous Maxwell demon – We now use our thermodynamic framework to analyze the purification process, and show that it can be interpreted as an autonomous Maxwell demon. We assume that the qubit is initially prepared in the maximally mixed state $\rho_{Q}(0)=\mathds{1}/2$. Fig. 4 shows the evolution of the qubit and cavity entropies, together with their quantum mutual information $I_{CQ}(t)$. We also plotted the work $W_{C}(t)$ and heat $Q_{C}(t)$ provided by the cavity, together with its internal energy variation $\Delta E_{C}$. We first note that the machine is able to decrease the qubit entropy (initially equal to its maximal value $\log 2$) down to a minimum value reached at time $t_{\text{min}}$, at the expanse of work performed by the cavity, akin to a Szilard engine [38, 39].
Figure 4: (a) von Neumann entropy of the cavity $S_{C}$ and of the qubit $S_{Q}$, and mutual information between the qubit and cavity $I_{CQ}$. (b) Cavity heat $Q_{C}$, work $W_{C}(t)$ and internal energy variation $\Delta E_{C}(t)$. For both (a) and (b), $n_{0}=50$, $\bar{n}=1$. Plotted quantities are computed from exact calculation with truncated cavity Hilbert space.
We then build on the dynamical analysis above to interpret the evolution of the thermodynamic quantities associated with the purification process. As the qubit-cavity correlations associated with the measurement grow during time interval $[0,t_{c}]$, the mutual information increases from 00 (as expected for an initial factorized state) to almost its maximal value $\log 2$ reached around $t_{c}$. The reduced entropy of the cavity also increases by the same amount. This step corresponds to the transfer of information about the qubit into the demon’s memory. Over the interval $[t_{c},t_{\text{min}}]$, the autonomous feedback of the cavity induces conditional rotations of the qubit state in the equator of the Bloch sphere based on the qubit state. This process is able to decrease the qubit entropy, but has no work cost (as seen on Fig. 4(b)) as the qubit’s internal energy remains constant. This apparent violation of the second law of thermodynamics is enabled by the consumption of the mutual information built during the measurement stroke, as can be seen in Fig. 4(a). Consumption of mutual information for an effect analogous to work expenditure is precisely the signature of a Maxwell demon [40, 39, 41]. This behavior here emerges autonomously from the Jaynes-Cummings interaction, and is naturally captured by our thermodynamic framework. We finish by noting that the evolution over the time intervals $[t_{k},t_{k+1}]$, correspond to a new measurement step followed by another feedback step, with degraded performances due to the building up of intra-branch correlations.
In conclusion, we have analyzed the dynamics of a qubit coupled to a cavity prepared in a displaced coherent state of very large displacement amplitude, a state commonly considered as guaranteeing a classical drive behavior of the cavity. Owing to our framework for the thermodynamics of autonomous machines, we have shown that at short enough times, the cavity indeed behaves as an ideal work source inducing a time-dependent drive on the qubit. However, over a time-scale set by the qubit-cavity interaction strength, the cavity performs a measurement of the qubit state in a basis set by the initial cavity displacement phase, responsible for the collapse of the Rabi oscillations. We have shown that this measurement is followed by an autonomous feedback, a state-dependent drive able to purify the qubit. From the analysis of thermodynamic quantities, we relate this purification to the cavity behaving as an autonomous Maxwell demon, spending work to build mutual information with the qubit, and then consuming this mutual information decrease the qubit’s entropy. Those results demonstrate the potential of consistent thermodynamic frameworks, like the one of Ref. [30], to shed new light on complex quantum dynamics. The parameters needed to implement our scheme are within state-of-the-art, e.g. in superconducting qubit-cavity systems, where strong couplings with respect to damping rates can be achieved. Natural extensions of this work include optimizing the initial cavity state to improve the purification process, e.g. by attaining purer states at a fixed feedback time, and testing the robustness of the purified qubit state to cavity preparation and residual qubit-cavity detunings.
Acknowledgements – C.E. acknowledges funding from French National Research Agency (ANR) under project ANR-22-CPJ1-0029-01. We thank Alexia Auffèves for fruitful discussions.
Supplemental Material: The Jaynes Cummings model as an autonomous Maxwell demon
I A. Physical Model
We consider a qubit $\left(Q\right)$ interacting with a harmonic oscillator ($C$) through a resonant Jaynes-Cummings coupling. The total Hamiltonian is given by
$\hat{H}$ $=\hat{H}_{C}+\hat{H}_{Q}+\hat{V}_{QC}=\hbar\omega_{0}a^{\dagger}a+\hbar\omega_{0}\sigma_{+}\sigma_{-}+i\hbar g\left(a\sigma_{+}-a^{\dagger}\sigma_{-}\right)$ (S1)
with $\omega_{0}$ the common transition frequency and $g$ the coupling constant between the qubit and the cavity. The operators $a$ and $a^{\dagger}$ are the usual bosonic annihilation and creation operators and $\sigma_{+}=|e\rangle\langle g|=\sigma_{-}^{\dagger}$ with $|e\rangle$ and $|g\rangle$ the excited and ground states of the two-level system, respectively. The total unitary evolution operator can be written as follows [33]:
$\hat{U}\left(t\right)=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)|e\rangle\langle e|+\cos\left(gt\sqrt{a^{\dagger}a}\right)|g\rangle\langle g|+\frac{\sin\left({gt\sqrt{a^{\dagger}a+1}}\right)}{\sqrt{a^{\dagger}a+1}}a\sigma_{+}-\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}\sigma_{-}$ (S2)
II B. Quasi-ideal work source behavior of the cavity
To verify the quasi-ideal work source behavior of the cavity, we consider the following initial factorized state for the qubit and cavity
$\hat{\rho}\left(0\right)=|e\rangle\langle e|\otimes D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}\;.$ (S3)
Under the unitary evolution the state $\hat{\rho}(t)=\hat{U}(t)\hat{\rho}\left(0\right)\hat{U}^{\dagger}(t)$ takes the form
$\hat{\rho}\left(t\right)$ $=|e\rangle\langle e|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$ (S4)
-- $|e\rangle\langle g|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
$-|g\rangle\langle e|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$
$+|g\rangle\langle g|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}\;.$
Tracing over the cavity Hilbert space, the reduced qubit state is given by
$\hat{\rho}_{Q}\left(t\right)=\begin{bmatrix}P_{e}\left(t\right)&C_{eg}\left(t\right)\\ C_{eg}\left(t\right)&1-P_{e}\left(t\right)\end{bmatrix}$ (S5)
with population $P_{e}$ and coherence $C_{eg}$
$P_{e}\left(t\right)$ $=\sum_{n}\cos^{2}\left({gt\sqrt{n+1}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (S6)
$C_{eg}\left(t\right)$ $=-\sum_{n}\cos\left({gt\sqrt{n+2}}\right)\sin\left({gt\sqrt{n+1}}\right)\langle n+1|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle\;.$
For $n_{0}\gg 1$ the terms $\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ and $\langle n+1|D\left(\alpha\right)W_{\beta\left(0\right)}D^{\dagger}\left(\alpha\right)|n\rangle$ as a function of $n$ are close to Gaussian distributions peaked around $n_{0}$ with variance much smaller than the mean. This allows us to expand the trigonometric functions in $P_{e}\left(t\right)$ and $C_{eg}\left(t\right)$ around $n_{0}$. In the following, we parameterize the population and coherence of qubit in terms of rotation angle $\theta=2gt\sqrt{n_{0}}$ in the Bloch sphere.
For large $n_{0}$:
$P_{e}\left(\theta\right)$ $=\sum_{n}\cos^{2}\left({\frac{\theta}{2}\sqrt{1+\frac{n-n_{0}+1}{n_{0}}}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (S7)
$=\sum_{n}[\cos^{2}{\frac{\theta}{2}}-\frac{\theta\sin{\theta}}{4}\left(\frac{n-n_{0}+1}{n_{0}}\right)-\frac{\theta\left(\theta\cos{\theta}-\sin{\theta}\right)}{16}\left(\frac{n-n_{0}+1}{n_{0}}\right)^{2}+O\left(\frac{1}{n_{0}^{3}}\right)]$
×$\times\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$
$=\cos^{2}{\frac{\theta}{2}}-\frac{\theta\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)}{16n_{0}}-\frac{\theta\left(1+\bar{n}\right)\left(1+2\bar{n}\right)\left(\theta\cos{\theta}-\sin{\theta}\right)}{16n_{0}^{2}}+O\left(\frac{1}{n_{0}^{3}}\right)$
$=\cos^{2}{\frac{\theta}{2}}+\frac{\delta P_{e}\left(\theta,\bar{n}\right)}{n_{0}}+O\left(\frac{1}{n_{0}^{2}}\right)$
where
$\delta P_{e}\left(\theta,\bar{n}\right)=-\frac{\theta}{16}\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)\;.$ (S8)
In a similar way
$C_{eg}\left(\theta\right)$ $=-e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{16n_{0}}+$ (S9)
$+\frac{\left(5+8\bar{n}\right)2\theta-\left(11+4\bar{n}\left(11+9\bar{n}\right)\right)2\theta\cos{\theta}+\left(12+72\bar{n}\left(1+\bar{n}\right)-\left(9+8\bar{n}\left(4+3\bar{n}\right)\right)\theta^{2}\right)\sin{\theta}}{64n_{0}^{2}}$
$+O\left(\frac{1}{n_{0}^{3}}\right)]$
$=e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{\delta C_{eg}\left(\theta,\bar{n}\right)}{n_{0}}+O\left(\frac{1}{n_{0}^{2}}\right)]$
where
$\delta C_{eg}\left(\theta,\bar{n}\right)=\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{n_{0}}$ (S10)
We use the same methodology to calculate the expansion of the thermodynamic quantities of the cavity. To determine those thermodynamics quantities, we observe from the numerics that, in the unitary regime, the cavity remains approximately in a Gaussian state which is thus fully characterized by $\langle a\rangle$, $\langle a^{\dagger}\rangle$, $\langle a^{\dagger}a\rangle$ and $\langle a^{2}\rangle$.
Starting with the exact reduced state of the cavity for $\phi_{0}=0$
$\hat{\rho}_{C}\left(t\right)$ $=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)+$ (S11)
$\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
the average $\langle a\rangle$ is given by
$\langle a\rangle$ $=Tr[\hat{\rho}_{C}\left(t\right)a]$ (S12)
$=\sum_{n}[\cos\left(gt\sqrt{n+1}\right)\cos\left(gt\sqrt{n}\right)+{\sin\left({gt\sqrt{n+1}}\right)}\sin\left({gt\sqrt{n}}\right)\frac{\sqrt{n+1}}{\sqrt{n}}]\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle$
which can be decomposed into
$\langle a\rangle$ $=\sum_{n}\left(\Lambda_{0}+\Lambda_{1}n+\Lambda_{2}n^{2}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle+O\left(\frac{1}{n_{0}^{4}}\right)$ (S13)
with
$\Lambda_{0}=\frac{{1+\frac{\theta^{2}}{4}+2n_{0}\left(-3+4n_{0}\left(3+4n_{0}\right)-\theta^{2}\right)+\left(-1+6n_{0}-24n_{0}^{2}+\left(1-2n_{0}\right)^{2}\frac{\theta^{2}}{4}\right)\cos\left(\theta\right)-\left(1-3n_{0}\right)^{2}\theta\sin\left(\theta\right)}}{{32n_{0}^{3}}}$ (S14)
$\Lambda_{1}=\frac{{1-6n_{0}+\frac{\theta^{2}}{4}+\left(-1+\frac{\theta^{2}}{4}-2n_{0}\left(-3+\frac{\theta^{2}}{4}\right)\right)\cos\left(\theta\right)+\left(-2+7n_{0}\right)\frac{\theta}{2}\sin\left(\theta\right)}}{{8n_{0}^{3}}}$ (S15)
$\Lambda_{2}=\frac{{4+2\left(\frac{\theta^{2}}{4}-2\right)\cos\left(\theta\right)-\frac{5\theta}{2}\sin\left(\theta\right)}}{{16n_{0}^{3}}}$ (S16)
Summing up these terms we get
$\langle a\rangle$ $=\sqrt{n_{0}}\Big{(}1+\frac{\sin^{2}{\frac{\theta}{2}}}{2{n_{0}}}+\frac{-2\sin^{2}{\frac{\theta}{2}}\left(1+\frac{\theta^{2}}{2}\right)+\frac{\theta}{2}\sin{\theta}+\bar{n}\theta\left(\theta\cos{\theta}-\sin{\theta}\right)}{16{n_{0}}^{2}}+$ (S17)
$\frac{13+\frac{5\theta^{2}}{4}+8\bar{n}\left(7+6\bar{n}+\frac{\theta^{2}}{4}\right)+\left(-13+\frac{9\theta^{2}}{4}+8\bar{n}\left(-7-6\bar{n}+\left(4+3\bar{n}\right)\frac{\theta^{2}}{4}\right)\right)\cos{\theta}-2\left(5+\bar{n}\left(19+15\bar{n}\right)\right)\theta\sin{\theta}}{32n_{0}^{3}}$
$+O\left(\frac{1}{n_{0}^{4}}\right)\Big{)}\;.$
Similarly, we find
$\langle a^{2}\rangle$ $=n_{0}\Big{(}1+\frac{\sin^{2}{\frac{\theta}{2}}}{{n_{0}}}+\frac{\left(1+2\bar{n}\right)[\left(1+\frac{\theta^{2}}{4}\right)\cos{\theta}-1]+\left(3+2\bar{n}\right)\frac{\theta}{4}\sin{\theta}-\frac{\theta^{2}}{2}}{4{n_{0}}^{2}}$ (S18)
$+\frac{5+\frac{3\theta^{2}}{4}+6\bar{n}\left(4+4\bar{n}+\frac{\theta^{2}}{4}\right)+\left(-5+\theta^{2}+3\bar{n}\left(-8\left(1+\bar{n}\right)+\left(5+4\bar{n}\right)\frac{\theta^{2}}{4}\right)\right)\cos\theta-\left(17+69\bar{n}+60\bar{n}^{2}\right)\frac{\theta}{4}\sin{\theta}}{4n_{0}^{3}}$
$+O\left(\frac{1}{n_{0}^{4}}\right)\Big{)}\;.$
Owing to the resonance condition, the total Hamiltonian $\hat{H}$ conserves the total number of excitations. Consequently, we have at any time:
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+1-P_{e}\left(t\right)$ (S19)
Using the expansion of $P_{e}\left(t\right)$ up to second order Eq. (S7), we simplify the above term as
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+\sin^{2}{\frac{\theta}{2}}+\frac{\theta\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)}{16n_{0}}+\frac{\theta\left(1+\bar{n}\right)\left(1+2\bar{n}\right)\left(\theta\cos{\theta}-\sin{\theta}\right)}{16n_{0}^{2}}$ (S20)
$+O\left(\frac{1}{n_{0}^{3}}\right)\;.$
The covariance matrix is defined as:
$V$ $=\begin{bmatrix}\sigma_{q}^{2}&\sigma_{qp}\\ \sigma_{pq}&\sigma_{p}^{2}\\ \end{bmatrix}$ (S21)
with
[…]
Since $\phi_{0}=0$, the covariance matrix $V$ takes the diagonal form
$V$ $=\begin{bmatrix}\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}-2\langle a\rangle^{2}&0\\ 0&-\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}\\ \end{bmatrix}\;.$ (S22)
For Gaussian states, the von Neumann entropy can be evaluated as
$S_{C}\left(t\right)$ $=h\left(\sqrt{Det[V]}\right),\hskip 5.0pth\left(x\right)=\left(x+\frac{1}{2}\right)\log\left({x+\frac{1}{2}}\right)-\left(x-\frac{1}{2}\right)\log\left({x-\frac{1}{2}}\right)\;.$ (S23)
Using the previous expansions the determinant is expressed as
$\sqrt{Det[V]}$ $=\left(\bar{n}+\frac{1}{2}\right)\sqrt{[1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)]}$ (S24)
Using the framework of [30] we can calculate the heat exchanges of the cavity. We must first find the thermal state $\hat{\rho}_{C}^{th}\left(t\right)$ with the same von Neumann entropy $S_{C}\left(t\right)$ to do so. The von Neumann entropy of a thermal state with photon numbers $n_{c}$ is given by $\log[\frac{n_{c}^{n_{c}}}{\left(n_{c}+1\right)^{n_{c}+1}}]$. Comparing the entropies
$\log[\frac{n_{c}^{n_{c}}}{\left(n_{c}+1\right)^{n_{c}+1}}]=S_{C}\left(t\right)\implies n_{c}+\frac{1}{2}=\sqrt{Det[V]}$
To find the heat exchanges $Q_{C}$, we recall that the heat is defined as the negative of the variation in the thermal energy of the cavity, which takes the following form $Q_{C}\left(t\right)=-Tr[\hat{H}_{C}\left(\hat{\rho}_{C}^{th}\left(t\right)-\hat{\rho}_{C}^{th}\left(0\right)\right)]$ and reads
$Q_{C}\left(t\right)=\hbar\omega_{0}\left(\bar{n}+\frac{1}{2}\right)\left[1-\sqrt{1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)}\right]\;.$ (S25)
By using the expression for $\langle a^{\dagger}a\rangle$, we can also determine the change in the cavity’s internal energy as follows
$\Delta E_{C}\left(t\right)=\hbar\omega_{0}\left(1-P_{e}\left(t\right)\right)$ (S26)
In the end, the work done by the cavity is given by
$W_{C}\left(t\right)=-\Delta E_{C}\left(t\right)-Q_{C}\left(t\right)\;.$ (S27)
III C. Joint atom-cavity evolution on longer times
III.1 Measurement by the cavity
In this section, we provide the expressions for cavity and atom, evolution over the interval $[0,t_{c}]$, when the system starts in state:
$\hat{\rho}\left(0\right)=\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{C}\left(0\right),\hskip 5.0pt\hat{\rho}_{C}\left(0\right)=D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}$ (S28)
where in $\{|e\rangle,|g\rangle\}$ basis:
|±$|\pm_{y}\rangle\langle\pm_{y}|=\frac{1}{2}\begin{bmatrix}1&\mp i\\ \pm i&1\end{bmatrix},\hskip 10.0pt|\pm_{y}\rangle\langle\mp_{y}|=\frac{1}{2}\begin{bmatrix}1&\pm i\\ \pm i&-1\end{bmatrix}.$ (S29)
It is useful to recast the total evolution operator as
$\hat{U}\left(t\right)=\hat{c}_{1}|e\rangle\langle e|+\hat{c}_{0}|g\rangle\langle g|+\hat{s}_{1}|e\rangle\langle g|-\hat{s}_{0}|g\rangle\langle e|,$ (S30)
with $\hat{c}_{1}=\cos\left(gt\sqrt{a^{\dagger}a+1}\right),\hskip 1.0pt\hat{c}_{0}=\cos\left(gt\sqrt{a^{\dagger}a}\right),\hskip 1.0pt\hat{s}_{1}=\frac{\sin\left(gt\sqrt{a^{\dagger}a+1}\right)}{\sqrt{a^{\dagger}a+1}}a,\hskip 1.0pt\hat{s}_{0}=\frac{\sin\left(gt\sqrt{a^{\dagger}a}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}$ and to introduce
$\hat{s}_{1}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}},\hskip 5.0pt\hat{s}_{0}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (S31)
Below, we compute separately the evolution of the basis of qubit operators $\{|\nu_{y}\rangle\langle\mu_{y}|\}$, $\mu,\nu=\pm$.
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\hat{U}\left(t\right)\frac{1}{2}[|e\rangle\langle e|-i|e\rangle\langle g|+i|g\rangle\langle e|+|g\rangle\langle g|]\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ (S32)
$=\sum_{p,q=e,g}|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$
$=\sum_{p,q=e,g}\hat{T}_{pq},$
where $\hat{T}_{pq}=|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$. For $p=e$ and $q=e$:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{1}^{\dagger}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{1}^{\dagger}+i\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}].$ (S33)
We can use the definition of $\hat{s}^{{}^{\prime}}_{0,1}$ along with the relations $D^{\dagger}\left(\alpha_{0}\right)aD\left(\alpha_{0}\right)=a+\alpha$ and $D^{\dagger}\left(\alpha_{0}\right)a^{\dagger}D\left(\alpha_{0}\right)=a^{\dagger}+\alpha^{*}$ to transform the above into the following
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{C}\left(0\right)a^{\dagger}\hat{s}_{1}^{{}^{\prime}}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)a^{\dagger}\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{C}\left(0\right)\hat{c}_{1}]$ (S34)
$=\frac{1}{2}|e\rangle\langle e|\otimes\Big{[}\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+$
$\hat{s}_{1}^{{}^{\prime}}[n_{0}\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)]\hat{s}_{1}^{{}^{\prime}}$
$-i\hat{c}_{1}\left[\alpha^{*}\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\right]\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}[\alpha\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)]\hat{c}_{1}\Big{]}$
The terms that include at least one bosonic operator acting on the thermal state, such as $D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha$, $D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}$, and $D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)$, have a vanishing magnitude when tracing over the cavity, and are therefore neglected, leading to:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\alpha\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{1}-i\alpha^{*}\hat{s}_{1}^{{}^{\prime}}\right)$ (S35)
$=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{1}-i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)$
We can now use the following approximation, applicable for large $n_{0}$ to simplify the expression above:
$\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}+1}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}}$ (S36)
$=\cos\left(gt\sqrt{\hat{n}+1}\right)+i\sin\left(gt\sqrt{\hat{n}+1}\right)[1-\frac{1}{2}\frac{\hat{n}+1-n_{0}}{n_{0}}+..]$
$=e^{igt\sqrt{\hat{n}+1}}+O\left(\frac{1}{n_{0}}\right)$
$=e^{\frac{igt\sqrt{n_{0}}}{2}}e^{\frac{igt\left(\hat{n}+1\right)}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right).$
We obtain:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right).$ (S37)
We proceed similarly to simplify the other terms $\hat{T}_{pq}$. For $p=e$ and $q=g$:
$\hat{T}_{eg}=\hat{T}_{ge}^{\dagger}$ $=\frac{1}{2}|e\rangle\langle g|\otimes[-\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{0}^{\dagger}+\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{0}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{0}-i\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{0}^{\dagger}]$ (S38)
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\hat{s}_{1}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{0}-i\hat{s}_{0}^{\dagger}\right)$
$=\frac{-i}{2}|e\rangle\langle g|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{0}-i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}\right)$
$=-\frac{ie^{\frac{igt}{2\sqrt{n_{0}}}}}{2}|e\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right)$
We approximated term $\hat{c}_{0}-i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}$ using:
$\hat{c}_{0}+i\sqrt{n_{0}}\hat{s}_{0}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (S39)
$=\cos\left(gt\sqrt{\hat{n}}\right)+i\sin\left(gt\sqrt{\hat{n}}\right)[1-\frac{1}{2}\frac{\hat{n}-n_{0}}{n_{0}}+..]$
$=e^{igt\sqrt{\hat{n}}}+O\left(\frac{1}{n_{0}}\right)$
$=e^{\frac{igt\sqrt{n_{0}}}{2}}e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right)$
Finally, for $p=g$ and $q=g$
$\hat{T}_{gg}$ $=\frac{1}{2}|g\rangle\langle g|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+{\cal O}\left(\frac{1}{n_{0}}\right)$ (S40)
Collecting the four contributions:
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+{\cal O}\left(\frac{1}{n_{0}}\right)$ (S41)
Using the same method, we compute the evolution of the three other basis elements:
$\hat{U}\left(t\right)|-_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{1}{2}\left[\begin{matrix}1&ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S42)
$\hat{U}\left(t\right)|+_{y}\rangle\langle-_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S43)
$\hat{U}\left(t\right)|-_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\frac{e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$ (S44)
We deduce the evolution of the system starting from an arbitrary initial qubit state:
$\hat{\rho}_{QC}\left(t\right)$ $=$ $\hat{U}\left(t\right)\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes D\left(\sqrt{n_{0}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}\right)\hat{U}^{\dagger}\left(t\right)$ (S47)
$=$ $\frac{\rho_{++}}{2}\left[\begin{matrix}1&-ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{-\frac{igt}{2\sqrt{n_{0}}}}&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{--}}{2}\left[\begin{matrix}1&ie^{\frac{-igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{igt}{2\sqrt{n_{0}}}}&&1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)+$
$+\frac{\rho_{+-}e^{igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{\frac{igt}{\sqrt{n_{0}}}}&ie^{\frac{igt}{2\sqrt{n_{0}}}}\\ ie^{\frac{igt}{2\sqrt{n_{0}}}}&&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)$
$+\frac{\rho_{-+}e^{-igt\sqrt{n_{0}}}}{2}\left[\begin{matrix}e^{-\frac{igt}{\sqrt{n_{0}}}}&-ie^{-\frac{igt}{2\sqrt{n_{0}}}}\\ -ie^{\frac{-igt}{2\sqrt{n_{0}}}}&-1\end{matrix}\right]\otimes D\left(\alpha_{0}e^{-\frac{igt}{2\sqrt{n_{0}}}}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}e^{\frac{igt}{2\sqrt{n_{0}}}}\right)+O\left(\frac{1}{n_{0}}\right)$
$=$ $\rho_{++}|+_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{++}\left(t\right)+\rho_{--}|-_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{--}\left(t\right)$
$+\rho_{+-}e^{igt\sqrt{n_{0}}+\frac{igt}{\sqrt{n_{0}}}}|+_{y}\left(t\right)\rangle\langle-_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{+-}\left(t\right)++\rho_{-+}e^{-igt\sqrt{n_{0}}-\frac{igt}{\sqrt{n_{0}}}}|-_{y}\left(t\right)\rangle\langle+_{y}\left(t\right)|\otimes\hat{\rho}_{C}^{-+}\left(t\right)$
$+O\left(\frac{1}{n_{0}}\right),$
where |±$|\pm_{y}\rangle\left(t\right)=\frac{1}{\sqrt{2}}\left(\begin{matrix}1\\ \pm ie^{\mp\frac{igt}{2\sqrt{n_{0}}}}\end{matrix}\right)$, and $\hat{\rho}_{C}^{\nu\mu}\left(t\right)$ is defined in Eq. (2) of main text.
We now show that the trace of the terms $\hat{\rho}_{C}^{\pm\mp}\left(t_{c}\right)$ becomes negligible at $t_{c}$. We have:
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\text{Tr}[D\left(\sqrt{n_{0}}e^{\pm i\frac{gt}{2\sqrt{n_{0}}}}\right)w_{\beta\left(0\right)}D^{\dagger}\left(\sqrt{n_{0}}e^{\mp i\frac{gt}{2\sqrt{n_{0}}}}\right)]$ (S48)
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin\frac{gt}{\sqrt{n_{0}}}}\text{Tr}[w_{\beta\left(0\right)}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)]$
$=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}\left(1-e^{-\beta\left(0\right)}\right)e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}\frac{1}{\pi}\int d^{\left(2\right)}\alpha\langle\alpha|e^{-\beta\left(0\right)a^{\dagger}a}D\left(\pm 2i\sqrt{n_{0}}\sin\left(\frac{gt}{2\sqrt{n_{0}}}\right)\right)|\alpha\rangle$
On using the known properties of coherent states and displacement operator to simplify the integrand above and carrying out the Gaussian integrals, we obtain
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm igt(\sqrt{n_{0}}+1/\sqrt{n_{0}})}e^{\pm in_{0}\sin{\frac{gt}{\sqrt{n_{0}}}}}e^{-2n_{0}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}e^{-\frac{4n_{0}\bar{n}}{\left(1+\bar{n}\right)^{2}}\sin^{2}{\frac{gt}{2\sqrt{n_{0}}}}}$ (S49)
As $n_{0}\to\infty$
$\text{Tr}[\hat{\rho}_{C}^{\pm\mp}\left(t\right)]$ $=e^{\pm 2igt\sqrt{n_{0}}}e^{-\frac{g^{2}t^{2}}{2}}e^{-\frac{g^{2}t^{2}\bar{n}}{\left(1+\bar{n}\right)^{2}}+\frac{\bar{n}}{\left(1+\bar{n}\right)^{2}}\frac{g^{4}t^{4}}{12n_{0}}},$ (S50)
which tends to 00.
Consequently, we can express the reduced state of the atom at time $t_{c}$ as:
$\hat{\rho}_{Q}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\left(t_{c}\right)\rangle\langle+_{y}\left(t_{c}\right)|+\rho_{--}|-_{y}\left(t_{c}\right)\rangle\langle-_{y}\left(t_{c}\right)|+O\left(\frac{1}{n_{0}}\right)$ (S51)
We can see that for times $t\in[0,g^{-1}]$, the term $e^{\pm\frac{igt}{2\sqrt{n_{0}}}}$ approaches 1 as $n_{0}$ approaches infinity, which allows us to rewrite the state of the qubit at $t_{c}$ as:
$\hat{\rho}_{Q}\left(t_{c}\right)$ $=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+O\left(\frac{1}{n_{0}}\right)$ (S52)
III.2 Autonomous feedback
On a longer timescale of order $t\sim\sqrt{n_{0}}\pi g^{-1}$, the cavity exerts an effective drive on the qubit that is dependent on the measurement result. Here we analyze the total dynamics along two branches associated with the qubit found in |±$|\pm_{y}\rangle$. As we are in the interaction picture, the total system evolves as follows:
$i\hbar\frac{d\hat{\rho}_{QC}(t)}{dt}$ $=\left[\hat{V}_{QC},\hat{\rho}_{QC}(t)\right].$ (S53)
Since the qubit-cavity state is factorized within each branch up to an order of $O\left(\frac{1}{n_{0}}\right)$, we can trace over the cavity space and get the effective evolution for qubit. Denoting $\hat{\rho}_{Q}^{\pm\pm}(t)$ the qubit states along each branch, with $\hat{\rho}_{Q}^{\pm\pm}(0)=|\pm_{y}\rangle\langle\pm_{y}|$, we have:
$i\hbar\frac{d}{dt}\hat{\rho}_{Q}^{\pm\pm}(t)$ $=\left[i\hbar g(\alpha_{\pm}(t)\sigma_{+}-\alpha_{\pm}^{*}(t)\sigma_{-}),\hat{\rho}_{Q}^{\pm\pm}(t)\right]$ (S54)
From Eq. (S54), we obtain the effective Hamiltonian acting on qubit
$\hat{H}_{Q}^{\text{eff}(\nu)}(t)$ $=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-}).$ (S55)